# Watermarking Algorithm for Large Language Models

### v1

@@ Appendix B Examples of Watermarked Text
Figures 4 and 5 provide examples of text generated with and without the SimMark watermark. These examples illustrate the imperceptibility of the watermark to human readers while enabling robust detection through our proposed algorithm. They also highlight SimMark’s robustness to paraphrasing while maintaining quality comparable to non-watermarked text.
@@
Figure 4: Example of text generated with and without cosine-SimMark using the RealNews dataset. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=0.14\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;5.033$, false negative). The middle panel demonstrates text generated using SimMark with cosine similarity-based watermarking, producing a strong detection signal ($z_{\text{soft}}=9.48\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=6.94\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;5.033$).
@@
Figure 5: Example of text generated with and without Euclidean-SimMark using the BookSum dataset. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows the non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=-1.07\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@cmyk@stroke{0}{1}{1}{0}\pgfsys@color@cmyk@fill{0}{1}{1}{0}<}\;4.13$, false negative). The middle panel demonstrates text generated using SimMark with Euclidean distance-based watermarking, producing a strong detection signal ($z_{\text{soft}}=13.07\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite semantic alterations ($z_{\text{soft}}=11.99\;{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}>}\;4.13$).
Appendix C Ablation Study on Soft Count Smoothness Factor $K$
@@
[…]
@@
Appendix D Ablation Study on Impact of PCA
Table 4 presents the results of an ablation study investigating the impact of applying PCA to reduce the dimensionality of sentence embeddings across the RealNews, BookSum, and Reddit-TIFU datasets. Metrics include ROC-AUC, and TP at fixed FP rates (FP=1% and FP=5%). Higher values indicate better performance across all metrics, with PCA applied to embeddings to explore its effect on detection accuracy and robustness.
The results reveal that the effect of PCA depends on the choice of similarity measure. For Euclidean distance-based SimMark, applying PCA generally improves robustness against paraphrasing attacks across most datasets, except for the BookSum dataset. This improvement likely arises because reducing dimensionality helps mitigate noise in the embeddings, especially after paraphrase attack.
On the other hand, for cosine similarity-based SimMark, applying PCA reduces performance across all datasets. This reduction may be due to PCA altering the embeddings in a way that disrupts the angular relationships critical for cosine similarity calculations. These findings highlight the importance of adapting PCA usage based on the similarity measure employed to achieve optimal watermarking performance.
@@ Appendix D Ablation Study on Impact of PCA
RealNews Cosine-SimMark (No PCA) 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2
@@
BookSum Cosine-SimMark (No PCA) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5
@@
Reddit-TIFU Cosine-SimMark (No PCA) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4
@@
Appendix E Finding an Optimal Interval
Figure 6 shows the distribution of distances between embeddings of consecutive sentences for both human and LLM-generated text, calculated on a sample of size 1000 from the BookSum dataset (no PCA applied to the embeddings in this case). A small but noticeable distribution shift between the two can be observed. Based on this, the interval $[0.4,0.55]$ appears to be a reasonable choice for SimMark watermarking in this case.
It is important to note that changes to the embedding representations, such as applying PCA or using a different embedding model, will lead to altered distance distributions. Consequently, the interval must be adjusted accordingly to maintain optimal performance. For instance, if PCA is applied, the interval $[0.28,0.36]$ is suitable. Similarly, if we plot the figure for when cosine similarity is used instead of Euclidean distance, intervals $[0.81,0.94]$ and $[0.68,0.76]$ are good candidates for cases with and without PCA, respectively. This variability in the distance distributions also strengthens the algorithm’s resistance to reverse engineering.
Figure 6: Distribution of Euclidean distances between embeddings of consecutive sentences for human-written and LLM-generated text on the BookSum dataset, generated using OPT-1.3B. The figure demonstrates that the interval $[0.4,0.55]$ is a reasonable choice for Euclidean-SimMark in this case, though it is not necessarily the only viable option.
Appendix F Computing Threshold $\beta$ for soft-$z$-test
Recall that a text is classified as LLM-generated when $z_{\mathrm{soft}}>\beta,$ and as human-written otherwise. $z_{\mathrm{soft}}$ is the statistic used in the statistical test described in Eq. (2). To determine the threshold $\beta$ that limits the FP rate to 1% or 5%, we first need to estimate $p_{0}$, the probability that the consecutive embeddings’ similarity or distance falls within the predefined interval $[a,b]$. This value of $p_{0}$ is a key component in calculating the $z_{\mathrm{soft}}$, as it represents the proportion of valid sentences in human-written text under the given interval. $p_{0}$ serves as an indicator of how frequently valid sentences are expected to occur in human-written text.
To compute $p_{0}$, we analyze the distribution of similarities (or distances) using a histogram approach, such as the one depicted in Figure 6. Specifically, we employ a binning technique to approximate the area under the curve of distribution in the interval $[a,b]$. The process involves dividing the entire range of distances or similarities into a fixed number of bins—1000 bins in our implementation. Each bin represents a small segment of the range, and the histogram is used to calculate the proportion of samples that fall within the interval $[a,b]$.
Mathematically, $p_{0}$ is estimated as:
$\small p_{0}=\frac{\text{Number of samples in bins corresponding to }[a,b]}{\text{Size of the dataset}},$
@@ Appendix F Computing Threshold $\beta$ for soft-$z$-test
It is worth noting that the distribution of similarities or distances may vary depending on different factors such as embedding model, and similarity measure (e.g., cosine or Euclidean). As a result, $p_{0}$ and therefore $\beta$ are determined programmatically during the detection to ensure reliable performance of the watermarking algorithm.
Appendix G Analysis of Sampling Efficiency
The average number of samples required to generate a valid sentence is influenced by the chosen interval $[a,b]$. To provide further insights into this relationship, we estimated the area under the curve (AUC) of the embedding similarity distribution for an unwatermarked LLM. Specifically, for the interval $[0.68,0.76]$ (for cosine-SimMark), the estimated AUC is approximately $0.194$.
Using the mean of the geometric distribution, which is given by $\frac{1}{p}$, where $p$ is the probability of success (in this case, the probability of falling within the interval), this translates to an expected average of $\frac{1}{0.194}\approx 5.1$ samples per valid sentence. This estimate is on par with the experimental results observed in our study. The AUC estimates were computed using the binning technique, as described in Appendix F.
This analysis underscores the importance of carefully selecting the interval $[a,b]$, as narrower intervals may increase the number of required samples, leading to reduced sampling efficiency but better performance, while broader intervals may compromise the effectiveness of the watermark. By understanding the interplay between the interval choice and sampling efficiency, we can better optimize SimMark’s performance for different embedding models.
Appendix H Experimental Environment
The majority of the experiments, including text generation and detection tasks, were conducted on a workstation equipped with an Intel Core i9 processor, 64GB of RAM, and an Nvidia RTX 3090 GPU with 24GB of VRAM. Some of the experiments involving bigram paraphrasing were performed on compute nodes with Nvidia V100 GPU with 32GB of VRAM.
Appendix I Prompts Used with GPT-3.5-Turbo
Table 5 presents the prompts we used to obtain paraphrases using GPT-3.5-Turbo (accessed via OpenAI API [^8] https://platform.openai.com/docs/api-reference) for both regular paraphrasing and more aggressive bigram paraphrasing attacks. By using the same prompts as Hou et al. (2023), we ensured that our results were directly comparable to those extracted from their paper, maintaining consistency in evaluation methodology.
@@ Appendix H Experimental Environment
[…]
@@
[…]
Table 5: Prompts used to generate paraphrases with GPT-3.5-Turbo for regular and bigram attacks. These are the same prompts used by Hou et al. (2023) for consistent and comparable evaluation. Here, sent represents the target sentence to rephrase, context includes all preceding sentences, and num_beams specifies the number of paraphrases generated for the bigram attack. A higher num_beams value indicates a more aggressive attack.

### v2

@@ Appendix H Examples of Watermarked Text
Figures 7 and 8 provide examples of text generated with and without the SimMark watermark using OPT-1.3B. These examples illustrate the imperceptibility of the watermark to human readers while enabling robust detection through our proposed algorithm. They also highlight SimMark’s robustness to paraphrasing while maintaining quality comparable to non-watermarked text.
@@
Figure 7: Example of text generated with and without cosine-SimMark using RealNews dataset and OPT-1.3B model. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=0.14\;{\color[rgb]{0,0,0}<}\;5.033$, false negative). The middle panel demonstrates text generated using SimMark with cosine similarity-based watermarking, producing a strong detection signal ($z_{\text{soft}}=9.48\;{\color[rgb]{0,0.5,0}>}\;5.033$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite paraphrasing ($z_{\text{soft}}=6.94\;{\color[rgb]{0,0.5,0}>}\;5.033$).
@@
Figure 8: Example of text generated with and without Euclidean-SimMark using BookSum dataset and OPT-1.3B model. The first sentence (in black) is the prompt for the model, the green sentences are valid, and red sentences are invalid/partially valid. Numbers in parentheses represent the soft count for partially valid sentences. The top panel shows the non-watermarked text, which fails to produce a significant detection signal ($z_{\text{soft}}=-1.07\;{\color[rgb]{0,0,0}<}\;4.13$, false negative). The middle panel demonstrates text generated using SimMark with Euclidean distance-based watermarking, producing a strong detection signal ($z_{\text{soft}}=13.07\;{\color[rgb]{0,0.5,0}>}\;4.13$). The bottom panel shows paraphrased watermarked text using GPT-3.5-Turbo, where the embedded watermark remains detectable despite paraphrasing ($z_{\text{soft}}=11.99\;{\color[rgb]{0,0.5,0}>}\;4.13$).
Appendix I Ablation Study on Soft Count Smoothness Factor $K$
@@
Table 6: Ablation study on the smoothness factor $K$ in soft counting (Eq. (1)) using the RealNews dataset, with Pegasus as the paraphraser. Metrics reported include ROC-AUC $\uparrow$, TP@FP=1% $\uparrow$, and TP@FP=5% $\uparrow$, from left to right. The last row ($K=\infty$) corresponds to regular counting with a step function in the interval $[a,b]$. A smoothness factor of $K=250$ provides a good balance between performance before and after paraphrase attacks for both cosine-SimMark and Euclidean-SimMark. Notably, while soft counting slightly reduces performance in the absence of paraphrasing, it demonstrates enhanced robustness against paraphrasing, yielding an increase across all metrics for Pegasus paraphraser and potentially larger gains against more advanced paraphrasers.
@@
Appendix J Ablation Study on Impact of PCA
Table 7 presents the results of an ablation study investigating the impact of applying PCA to reduce the dimensionality of sentence embeddings across RealNews, BookSum, and Reddit-TIFU datasets. Metrics include ROC-AUC, and TP at fixed FP rates (FP=1% and FP=5%). Higher values indicate better performance across all metrics, with PCA applied to embeddings to explore its effect on detection accuracy and robustness.
The results reveal that the effect of PCA depends on the choice of similarity measure. For Euclidean distance-based SimMark, applying PCA generally improves robustness against paraphrasing attacks across most datasets, except for the BookSum dataset. This improvement likely arises because reducing dimensionality helps mitigate noise in the embeddings, especially after the paraphrasing attack. On the other hand, for cosine similarity-based SimMark, applying PCA reduces performance across all datasets. This reduction may be due to PCA altering the embeddings in a way that disrupts the angular relationships critical for cosine similarity calculations. These findings highlight the importance of adapting PCA usage based on the similarity measure employed to achieve optimal watermarking performance.
@@
RealNews
Cosine-SimMark (No PCA) 99.7 / 96.9 / 98.8 99.2 / 90.3 / 98.2
@@
BookSum
Cosine-SimMark (No PCA) 99.8 / 98.8 / 99.5 99.5 / 93.3 / 98.5
@@
Reddit-TIFU
Cosine-SimMark (No PCA) 99.1 / 96.3 / 97.6 98.9 / 94.5 / 96.4
@@
Table 7: Ablation study on the impact of applying PCA to embeddings across three datasets. Metrics reported include ROC-AUC $\uparrow$, TP@FP=1% $\uparrow$, and TP@FP=5% $\uparrow$, respectively, from left to right. Higher values indicate better performance across all metrics. Bold and underlined numbers denote the highest and second-highest values, respectively. For cosine-SimMark, not applying PCA yields better results, while for Euclidean-SimMark, applying PCA improves performance except on the BookSum dataset.
Appendix K Finding an Optimal Interval
Figure 9 shows the distribution of distances between embeddings of consecutive sentences for both human and LLM-generated text, calculated on a sample of size 1000 from the BookSum dataset (no PCA applied to the embeddings in this case). A small but noticeable distribution shift between the two can be observed. Based on this, the interval $[0.4,0.55]$ appears to be a reasonable choice for SimMark in this case. It is important to note that changes to the embedding representations, such as applying PCA or using a different embedding model, will lead to altered distance distributions. Consequently, the interval must be adjusted accordingly to maintain optimal performance. For instance, if PCA is applied, the interval $[0.28,0.36]$ is suitable. Similarly, if we plot the figure for when cosine similarity is used instead of Euclidean distance, intervals $[0.81,0.94]$ and $[0.68,0.76]$ are good candidates for cases with and without PCA, respectively. This variability in the distance distributions may also strengthen the algorithm’s resistance to reverse engineering.
Figure 9: Distribution of Euclidean distances between embeddings of consecutive sentences for human-written and LLM-generated text on BookSum dataset, generated using OPT-1.3B. The figure demonstrates that the interval $[0.4,0.55]$ is a reasonable choice for Euclidean-SimMark in this case, though it is not necessarily the only viable option.
Appendix L Computing Threshold $\beta$ for soft-$z$-test
Recall that a text is classified as LLM-generated when $z_{\mathrm{soft}}>\beta$, and as human-written otherwise. $z_{\mathrm{soft}}$ is the statistic used in the statistical test described in Eq. (2). To determine the threshold $\beta$ that limits the FP rate to 1% or 5%, we first need to estimate $p_{0}$, the probability that the consecutive embeddings’ similarity or distance falls within the predefined interval $[a,b]$. This value of $p_{0}$ is a key component in calculating the $z_{\mathrm{soft}}$, as it represents the proportion of valid sentences in human-written text under the given interval. $p_{0}$ serves as an indicator of how frequently valid sentences are expected to occur in human-authored text.
To compute $p_{0}$, we analyze the distribution of similarities (or distances) using a histogram approach, such as the one depicted in Figure 9. Specifically, we employ a binning technique to approximate the area under the curve of distribution in the interval $[a,b]$. The process involves dividing the entire range of distances or similarities into a fixed number of bins—1000 bins in our implementation. Each bin represents a small segment of the range, and the histogram is used to calculate the proportion of samples that fall within the interval $[a,b]$. Mathematically, $p_{0}$ is estimated as:
@@
It is worth noting that the distribution of similarities or distances may vary depending on different factors such as the embedding model and similarity measure (e.g., cosine or Euclidean). As a result, $p_{0}$ and therefore $\beta$ are determined programmatically during the detection to ensure reliable performance of the watermarking algorithm.
Appendix M Human Evaluation of SimMark’s Imperceptibility
Figure 10: Screenshot of the instructions and consent form shown to participants before the human evaluation study. It describes the task, evaluation criteria, expected time commitment, and privacy assurances.
Human evaluation provides a more direct measure of imperceptibility for our watermarking algorithm. To assess this, we conducted a small-scale A/B test comparing outputs generated by SimMark to unwatermarked ones, using 10 randomly selected BookSum samples. Each sample pair consisted of one watermarked and one unwatermarked text.
Option Chosen Percentage
SimMark identified as watermarked 36.7%
No noticeable difference 23.3%
Unwatermarked identified as watermarked 40.0%
Table 8: Results of a small-scale human evaluation measuring the imperceptibility of SimMark. Participants were asked to identify which of two outputs (watermarked vs. unwatermarked) appeared to be watermarked across 10 random samples. Responses were nearly evenly split, indicating that SimMark watermarking is largely imperceptible to human readers.
Three volunteers (all master’s students in Canada, Iranian, aged between 18 and 25, with a Computer Science background) participated in the study. No sensitive data was collected beyond age range, academic level, and general field of study. Before starting the evaluation, participants were shown the consent and instruction form (Figure 10), which described the task, the absence of any risks, and the use of their responses solely for academic purposes. Participants provided explicit consent by filling out the form and then proceeded with the evaluation task.
After filling in the form, participants were asked to identify which option—A or B—was more likely to contain a watermark. We also included the “N” (no noticeable difference) option to avoid forcing a binary choice when participants genuinely could not distinguish between the texts, thereby making the evaluation more reliable given the scale of the evaluation.
@@
Previous ${context} \n$ Current sentence to paraphrase: {sent}
@@
Previous ${context} \n$ Paraphrase in {num_beams} different ways and return a numbered list: {sent}
Table 9: Prompts used to generate paraphrases with GPT-3.5-Turbo for regular and bigram attacks. These are the same prompts used by Hou et al. (2024a) for consistent and comparable evaluation. Here, “sent” represents the target sentence to rephrase, “context” includes all preceding sentences, and “num_beams” specifies the number of paraphrases generated for the bigram attack. A higher “num_beams” value indicates a more aggressive attack. Following Hou et al. (2024a), we set “num_beams” to 10 to have 10 rephrases of each sentence.
We focused our human evaluation on asking participants to detect watermarking, rather than assess fluency or preference. This directly targets SimMark’s design objective of imperceptibility: ensuring that even when readers are explicitly prompted to spot a watermark, they find it difficult to do so.
As shown in Table 8, on average, participants labeled the SimMark output as watermarked in 36.7% of cases, while in 40.0% of cases they mistakenly identified the unwatermarked text as watermarked. In 23.3% of comparisons, they reported no noticeable difference. These results are close to random guess, suggesting that SimMark’s watermark is largely imperceptible to human readers while maintaining natural fluency. A larger-scale evaluation is deferred to future work.
Appendix N Prompts Used with GPT-3.5-Turbo for Paraphrasing
Table 9 presents the prompts we used to obtain paraphrases using GPT-3.5-Turbo (accessed via OpenAI API [^2] https://platform.openai.com/docs/api-reference) for both regular paraphrasing and more aggressive bigram paraphrasing attacks [^3] Used “gpt-3.5-turbo-16k” model.. By using the same prompts as Hou et al. (2024a), we ensured that our results were directly comparable to those extracted from their paper, maintaining consistency in evaluation methodology.